# Federated LLMs Fine-tuned with Adaptive Importance-Aware LoRA

### v1

@@ Abstract
Federated fine-tuning of pre-trained Large Language Models (LLMs) enables task-specific adaptation across diverse datasets while preserving data privacy. However, the large model size and heterogeneity in client resources pose significant computational and communication challenges. To address these issues, in this paper, we propose a novel Heterogeneous Adaptive Federated Low-Rank Adaptation (LoRA) fine-tuned LLM framework (HAFL). To accommodate client resource heterogeneity, we first introduce an importance-based parameter truncation scheme, which allows clients to have different LoRA ranks, and smoothed sensitivity scores are used as importance indicators. Despite its flexibility, the truncation process may cause performance degradation. To tackle this problem, we develop an importance-based parameter freezing scheme. In this approach, both the cloud server and clients maintain the same LoRA rank, while clients selectively update only the most important decomposed LoRA rank-1 matrices, keeping the rest frozen. To mitigate the information dilution caused by the zero-padding aggregation method, we propose an adaptive aggregation approach that operates at the decomposed rank-1 matrix level. Experiments on the 20 News Group classification task show that our method converges quickly with low communication size, and avoids performance degradation when distributing models to clients compared to truncation-based heterogeneous LoRA rank scheme. Additionally, our adaptive aggregation method achieves faster convergence compared to the zero-padding approach.
@@
Large Language Model, Low-Rank Adaptation, Federated Learning
publicationid: pubid: 978-1-6654-3540-6/22/\$31.00 © 2025 Crown
@@ I Introduction
Large Language Models (LLMs) have exhibited exceptional performance in understanding and generating natural language across a wide range of tasks, including applications in chatbots and search engines[1]. To achieve optimal performance on specific tasks, these pre-trained models often require further fine-tuning. However, conventional fine-tuning methods are typically centralized, involving the collection of raw data on a single client, which raises significant privacy concerns. To address this issue, Federated Learning (FL) [2] has emerged as a promising solution by enabling collaborative model training across decentralized LLM agents, where only model weights are shared instead of raw data, thereby preserving data privacy. Integrating FL with LLMs allows for leveraging diverse data sources, resulting in more robust and generalized models.
@@
As FL increasingly relies on wireless networks to support collaboration across clients, communication has emerged as a critical bottleneck in fine-tuning LLMs. Unlike centralized settings, FL requires frequent transmission of model updates over limited-bandwidth connections, posing significant challenges for models with billions of parameters, such as LLaMA3-8B[3]. Even in half-precision, transmitting these parameters demands approximately 16 GB per update, resulting in delays that significantly impact training efficiency. Moreover, the large parameter size imposes high demands on the client’s computational capabilities.
@@
Parameter-Efficient Fine-Tuning (PEFT) has been proposed to alleviate the aforementioned challenges[4, 5]. One of the widely used PEFT methods is Low-Rank Adaptation (LoRA), which freezes the original pre-trained parameters of LLM and trains a smaller number of additional parameters instead[6]. This approach significantly reduces computational and storage requirements while maintaining model performance comparable to full fine-tuning. Instruction tuning, a technique to improve LLMs’ ability to follow diverse natural language instructions across various tasks, can benefit from such parameter-efficient methods. Authors in [7] proposed an approach that integrates FL with LoRA to enhance the instruction tuning of LLMs. This method enables clients to only train and optimize the LoRA layers using locally diverse instruction data, reducing privacy exposure risks by sharing only the LoRA weights, which significantly lowers the demands on communication and client computational capabilities. However, the study assumed a uniform LoRA rank across all clients and does not consider the potential impacts of client resource heterogeneity.
@@
To accommodate heterogeneous computational capacities, the authors in [8] proposed HETLoRA that assigned different LoRA ranks to different clients. Additionally, this work introduced a zero-padding aggregation method, and a global model truncation distribution method to address the challenges posed by distinct LoRA ranks. However, the zero-padding aggregation method can dilute the information learned by high-rank clients, and the truncation approach may lead to performance loss in the models distributed to clients.
@@
To address the challenges posed by different LoRA ranks, the authors in [9] synthesized a full-size LoRA through the direct multiplication of the A and B matrices, ensuring consistent dimensions of the full-size LoRA across clients. This full-size LoRA is then aggregated to form an updated global model, which is then decomposed using Singular Value Decomposition (SVD). The SVD components are further processed using a low-rank approximation to fit different client ranks before being distributed to the clients. However, the reconstruction method in [9] may lose information on the cross-relation across clients during aggregation, and the low-rank approximation distribution method still inevitably leads to performance loss.
In this paper, to address the limitations of previous work [7, 8, 9], we propose a novel Heterogeneous Adaptive Federated LoRA fine-tuned LLMs framework (HAFL) to handle the challenges posed by heterogeneous client resources. This framework ensures that each client’s LoRA rank is the same as the rank of the global LoRA at the cloud server for model compatibility. Each client selectively fine-tunes a portion of the LoRA layers based on its computational capabilities, keeping the remaining layers frozen rather than being truncated. Compared to HETLoRA using different ranks for different clients[8], our approach does not result in performance loss due to the truncation when adapting the highest-rank global LoRA to clients. Additionally, the freezing operation ensures that our scheme does not increase the consumption of training resources such as computational power and memory usage. Our contributions can be summarized as follows:
@@
  • We propose a federated LoRA fine-tuned LLM framework. To accommodate client resource heterogeneity, we first introduce an importance-based parameter truncation scheme, which allows clients to have different ranks to fine-tune the more critical LoRA layers. To avoid the performance degradation resulting from the truncation process, we further develop an importance-based parameter freezing scheme. In this approach, both the cloud server and clients maintain the same high LoRA rank, while clients selectively update only the most important LoRA rank-1 matrices while keeping others frozen.
@@
  • We propose an adaptive global model aggregation method for the cloud server. Rather than aggregating the entire LoRA matrices, we perform aggregation at the decomposed rank-1 matrix level. For each rank-1 matrix, only the clients that update and upload it will participate in the aggregation process. Unlike the zero-padding method, which allows clients with non-updated rank-1 matrices to join the aggregation process using zero values, our method effectively prevents information dilution.
  • We conduct experiments on the text classification task. The results show that our proposed HAFL method converges quickly with low communication overhead. Moreover, our importance-based partial freezing scheme effectively maintains client model accuracy during global model parameter distribution, outperforming the truncation approach. Additionally, our proposed adaptive aggregation method achieves faster convergence compared to the zero-padding approach in [8].
@@
The rest of the paper is organized as follows: Section II introduces the system model; Section III presents our novel importance-based parameters truncation and freezing schemes and our adaptive global aggregation; Section IV presents the numerical results; and finally, Section V concludes the paper.
Figure 1: HAFL Framework
@@ II System Model
As shown in Fig. 1, our proposed HAFL framework consists of a cloud server and a set of $N$ clients, denoted as $\mathcal{N}=\{1,2,\ldots,N\}$. Each client and the cloud server have a LoRA based LLM, where the parameters of the original pre-trained LLM are frozen, and only the incorporated low-rank decomposition matrices are trainable. Each client $k\in\mathcal{N}$ has a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$, where $x_{i}$ is the $i$th input data sample, $y_{i}$ is the corresponding labeled output, and $D_{k}$ is the number of data samples. Considering the limited communication resource, in each communication round, a subset of clients, $\mathcal{K}$ of size is randomly selected from the entire set of clients $\mathcal{N}$.
II-A Cloud LoRA Parameters Broadcast
The model updates of the LoRA based LLM at the cloud server and the clients can be expressed as
$\mathbf{W}=\mathbf{W}_{\text{pre}}+\Delta\mathbf{W}=\mathbf{W}_{\text{pre}}+\mathbf{B}\mathbf{A},$ (1)
where $\mathbf{W}_{\text{pre}}\in\mathbb{R}^{d\times l}$ denotes the fixed pre-trained model weights, and $\Delta\mathbf{W}\in\mathbb{R}^{d\times l}$ is the trainable matrix. The low-rank decomposition of $\Delta\mathbf{W}$ is given by $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times l}$ are the low-rank matrices with rank $r$. Here, $d$ and $l$ are the dimensions of the model’s weight matrix, and $r$ is the rank of the low-rank approximation, which is typically much smaller than both $d$ and $l$.
For different clients $k$, the LoRA rank is denoted as $r_{k}$. In our importance-based parameter truncation scheme, the client LoRA rank $r_{k}$ is selected from the range $[r_{\text{min}},r_{\text{max}}]$. In our importance-based parameter freezing scheme, the rank $r_{k}$ is set to be equal to $r_{\text{max}}$. Regardless of the scheme used, the LoRA rank of the global model at the cloud server is denoted as $r_{g}$, where $r_{g}=r_{\text{max}}$.
@@ II-B Local Model Training and Global Aggregation
After completing local training, clients upload their updated LoRA parameters $\{\mathbf{B}_{k}^{(t+1)},\mathbf{A}_{k}^{(t+1)}\}$ to the cloud server. Upon receiving updates from all clients, the cloud server aggregates these parameters to update the global LoRA parameters $\Theta_{\text{LoRA}}^{(t+1)}$, preparing for the next communication round. This approach allows the HAFL framework to effectively optimize the model using distributed data while preserving data privacy.
III Importance-based Parameters Truncation and Freezing Schemes

### v2

# HAFLQ: Heterogeneous Adaptive Federated LoRA Fine-tuned LLM with Quantization

@@ Abstract
Federated fine-tuning of pre-trained Large Language Models (LLMs) enables task-specific adaptation across diverse datasets while preserving privacy. However, challenges such as high computational and memory demands, heterogeneous client resources, bandwidth constraints, and ineffective global aggregation hinder its efficiency. To address these issues, we propose HAFLQ (Heterogeneous Adaptive Federated Low-Rank Adaptation Fine-tuned LLM with Quantization), a novel framework for efficient and scalable federated fine-tuning of LLMs in heterogeneous environments. To reduce memory and computation demands, we propose a salience-driven adaptive LLM quantization framework that evaluates the importance of transformer blocks using a salience metric and applies adaptive block-wise quantization accordingly. To handle heterogeneous computational capabilities, we propose an importance-based parameter truncation and freezing scheme. To address communication bottlenecks, we propose an importance-aware bandwidth-adaptive quantization method, which dynamically adjusts parameter precision based on importance and bandwidth constraints. To improve global model aggregation, we propose an adaptive rank-1 matrix-level aggregation strategy, which prevents information dilution and accelerates convergence by aggregating only updated rank-1 matrices from clients. Experimental results on the text classification task demonstrate that HAFLQ reduces memory usage by 31%, lowers communication cost by 49%, improves accuracy by 50%, and achieves faster convergence compared to the baseline method.
@@
Large Language Model, Low-Rank Adaptation, Federated Learning, Quantization.
@@ I Introduction
Large Language Models (LLMs) have exhibited exceptional performance in understanding and generating natural language across a wide range of tasks, including applications in chatbots and search engines[2, 3, 4]. To achieve optimal performance on specific tasks, these pre-trained models often require further fine-tuning. However, conventional fine-tuning methods are typically centralized, involving the collection of raw data on a single client, which raises significant privacy concerns. Federated Learning (FL), as a privacy-preserving distributed Machine Learning (ML) training paradigm[5, 6, 7], is a promising approach to address this issue by enabling collaborative fine-tuning across decentralized LLM agents without aggregating raw datasets. Integrating FL with LLMs not only addresses privacy concerns but also leverages diverse data sources from decentralized clients to fine-tune LLMs on specific domains.
@@
As FL increasingly relies on wireless networks to support collaboration across clients, communication has emerged as a critical bottleneck in fine-tuning LLMs. Unlike centralized settings, FL requires frequent transmission of model updates over limited-bandwidth connections, posing significant challenges for models with billions of parameters, such as LLaMA3-8B[8]. Even in half-precision, transmitting these parameters demands approximately 16 GB per update, resulting in delays that significantly impact training efficiency. Moreover, the large parameter size imposes high demands on the client’s computational capabilities. For instance, full fine-tuning of LLaMA3-8B requires memory for model parameters, optimizer states (twice the size of the parameters), and gradients (equivalent to the size of the parameters). Even with mixed precision, it requires at least 64 GB of GPU memory per client, posing significant challenges for resource-constrained devices.
To address these challenges, LLM quantization [9, 10, 11] has been proposed as a solution to alleviate computational resource constraints, such as GPU memory usage, in federated LLM fine-tuning. The authors in [12] proposed FedPipe, which applied low-precision quantization to reduce computational resource consumption. However, this approach employs a uniform quantization strategy across all clients, overlooking the potential heterogeneity in their computational capabilities. In practice, clients with powerful GPUs can support full-precision LLM fine-tuning without quantization, while aggressive quantization on such devices may lead to unnecessary performance loss. Conversely, resource-constrained clients with limited GPU memory or computational power can selectively apply higher levels of quantization to adapt to their resource limitations, highlighting the need for more adaptive strategies.
@@
In addition to quantization, Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed to further alleviate the computational and communication challenges in federated LLM fine-tuning [13, 14, 15]. One of the widely used PEFT methods is Low-Rank Adaptation (LoRA), which freezes the original pre-trained parameters of the LLM and trains a smaller number of additional parameters instead[16]. This approach significantly reduces computational and storage requirements while maintaining a model performance comparable to full fine-tuning. The authors in [17] proposed an approach that integrates FL with LoRA to enhance the instruction tuning of LLMs. This method enables clients to only train and optimize the LoRA layers using locally diverse instruction data, reducing privacy exposure risks by sharing only the LoRA weights, which significantly lowers the demands on communication and client computational capabilities. However, the study assumed a uniform LoRA rank across all clients and does not consider the potential impact of client resource heterogeneity.
@@
To accommodate heterogeneous computational capabilities, the authors in [18] proposed HETLoRA that assigned different LoRA ranks to different clients. To facilitate the aggregation and distribution of LoRA matrices with different ranks, this work introduced a zero-padding aggregation method and a global model truncation distribution method. However, the zero-padding aggregation method can dilute the information learned by high-rank clients, and the truncation approach may lead to performance loss in the models distributed to clients.
@@
To address the challenges posed by different LoRA ranks, the authors in [19] synthesized a full-size LoRA through the direct multiplication of the LoRA A and B matrices, ensuring consistent dimensions of the full-size LoRA across clients. This full-size LoRA is then aggregated to form an updated global model, which is then decomposed using Singular Value Decomposition (SVD). The SVD components are further processed using a low-rank approximation to fit different client ranks before being distributed to the clients. However, the reconstruction method in [19] may lose information regarding the cross-relation across clients during aggregation, and the low-rank approximation distribution method still inevitably leads to performance loss.
On the other hand, to further reduce communication costs in federated LoRA fine-tuned LLMs, the authors in [20] proposed integrating a quantization strategy for LoRA parameters to adapt to bandwidth-limited scenarios. However, this approach utilized fixed quantization precision across all clients, and failed to account for clients’ communication heterogeneity. For instance, clients with good channel conditions could support higher precision to accelerate convergence, whereas clients with poor channel conditions might require lower precision quantization to alleviate communication bottlenecks. Moreover, their methods did not consider the importance of individual rows and columns in the LoRA A and B matrices, treating all parameters equally during transmission. This lack of prioritization can lead to inefficient communication, as important components may be lost due to bandwidth constraints, potentially slowing down convergence and degrading global model performance.
To address the limitations of previous work [17, 18, 19, 20, 12], we propose a novel Heterogeneous Adaptive Federated LoRA Fine-tuned LLM with Quantization (HAFLQ) framework to handle the challenges posed by heterogeneous client computational and communication resources. Our contributions can be summarized as follows:
  • Salience-Driven Adaptive Quantization for Efficient LLM Deployment. To enable efficient deployment of LLMs on heterogeneous devices, we propose a salience-driven and resource-aware adaptive quantization scheme. By ranking transformer blocks based on their contribution to model performance, our method applies higher precision to more important blocks while quantizing less important ones. Additionally, the quantization strategy is dynamically adapted to the computational capabilities and memory constraints of each client device. Clients with more computational resources retain more blocks in full precision, while resource-constrained clients apply lower-precision quantization to a larger proportion of blocks.
@@
  • Importance-Aware LoRA Fine-Tuning with Truncation and Freezing. To accommodate client resource heterogeneity, we propose an importance-based parameter truncation scheme, which allows clients to fine-tune LoRA layers with different ranks based on their computational capabilities. To avoid performance degradation caused by truncation, we further develop an importance-based parameter freezing scheme. In this approach, both the cloud server and the clients maintain the same high LoRA rank, while clients selectively update only the most important LoRA rank-1 matrices while keeping others frozen.
  • Importance-Aware Bandwidth-Adaptive Communication Quantization. To address communication heterogeneity in wireless FL systems, we propose an Importance-Aware Bandwidth-Adaptive Communication Quantization scheme. This method dynamically adjusts the precision of the transmitted parameters based on their importance and each client’s bandwidth budget. By prioritizing the transmission of the most important LoRA rank-1 matrices with higher precision, the scheme optimizes bandwidth utilization and improves FL performance under constrained communication conditions.
@@
  • Adaptive Rank-1 Matrix-Level Model Aggregation. To improve global model aggregation, we propose an adaptive aggregation method at the rank-1 matrix level. Instead of aggregating entire LoRA matrices, only the clients that update and upload specific rank-1 matrices participate in the aggregation process. This prevents information dilution and accelerates convergence compared to traditional zero-padding methods.
  • We evaluate our methods on the Banking 77 text classification task. The salience-driven adaptive LLM quantization method reduces memory usage by 31% and computational costs, enabling efficient deployment on heterogeneous clients without sacrificing performance. Additionally, our importance-based partial freezing scheme preserves client model accuracy better than truncation-based methods, while the adaptive aggregation method achieves faster convergence compared to the zero-padding approach in [18]. Finally, our Importance-Aware Bandwidth-Adaptive Quantization optimizes bandwidth usage, reducing communication cost by 49%, while maintaining high accuracy under bandwidth constraints.
Figure 1: HAFLQ framework.
@@
The rest of the paper is organized as follows: Section II introduces the system model. Sections III and IV describe the salience-driven adaptive LLM quantization scheme and the importance-based parameter truncation and freezing schemes, respectively. Section V proposes an importance-aware and bandwidth-adaptive communication quantization scheme, followed by Section VI, which presents the adaptive global aggregation scheme. Numerical results are discussed in Section VII, and conclusions are summarized in Section VIII.
@@ II System Model
As shown in Fig. 1, our proposed HAFLQ framework consists of a cloud server, a base station, and a set of $K$ clients, denoted as $\mathcal{K}=\{1,2,\cdots,K\}$. Each client and the cloud server have a LoRA based LLM, where the parameters of the original pre-trained LLM are frozen, and only the incorporated low-rank decomposition matrices are trainable. Each client $k\in\mathcal{K}$ has a local dataset $\mathcal{D}_{k}=\{(x_{i},y_{i})\}_{i=1}^{D_{k}}$, where $x_{i}$ is the $i$th input data sample, $y_{i}$ is the corresponding labeled output, and $D_{k}$ is the number of data samples. Each client connects to the base station via wireless channels to transmit parameters to the cloud server.
II-A Cloud LoRA Parameter Broadcast
The updated model parameters of the LoRA-based LLM are handled differently at the cloud server and the clients to address their respective computational and memory constraints. Specifically, the cloud server retains the pre-trained model weights $\mathbf{W}_{\text{pre}}\in\mathbb{R}^{d\times l}$ in full precision, while the clients use a quantized version of the pre-trained weights, denoted as $\mathbf{W}_{\text{pre,q}}\in\mathbb{R}^{d\times l}$. For the cloud server, the model parameters can be expressed as
$\mathbf{W}_{\text{server}}=\mathbf{W}_{\text{pre}}+\Delta\mathbf{W}=\mathbf{W}_{\text{pre}}+\mathbf{B}\mathbf{A}.$ (1)
For the client, the model parameters can be expressed as
$\mathbf{W}_{\text{client}}=\mathbf{W}_{\text{pre,q}}+\Delta\mathbf{W}=\mathbf{W}_{\text{pre,q}}+\mathbf{B}\mathbf{A},$ (2)
where $\Delta\mathbf{W}\in\mathbb{R}^{d\times l}$ is the trainable matrix. The low-rank decomposition of $\Delta\mathbf{W}$ is given by $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times l}$ are low-rank matrices. Here, $d$ and $l$ are the dimensions of the model’s weight matrix, and $r$ is the rank of the low-rank approximation, which is typically much smaller than both $d$ and $l$.
For different clients $k$, the LoRA rank is denoted as $r_{k}$. In our importance-based parameter truncation scheme, the client LoRA rank $r_{k}$ is selected from the range $[r_{\text{min}},r_{\text{max}}]$. In our proposed importance-based parameter freezing scheme, rank $r_{k}$ is set to be equal to $r_{\text{max}}$. Regardless of the scheme used, the LoRA rank of the global model at the cloud server is denoted as $r_{g}$, where $r_{g}=r_{\text{max}}$.
@@ II-B Local Model Training and Global Aggregation
After completing local training, clients adaptively quantize their updated LoRA parameters, $\{\mathbf{B}_{k}^{(t+1)},\mathbf{A}_{k}^{(t+1)}\}$, into $\{\mathbf{B}_{k,\text{q}}^{(t+1)},\mathbf{A}_{k,\text{q}}^{(t+1)}\}$ before uploading them to the cloud server. Unlike traditional fixed or uniform quantization methods, our proposed importance-aware bandwidth-adaptive quantization adjusts the precision of each parameter based on its importance and the client’s bandwidth constraints. This ensures that the most critical rank-1 matrices are prioritized for transmission with higher precision, optimizing bandwidth utilization and minimizing performance degradation in heterogeneous environments.
Upon receiving updates, the cloud server dequantizes the parameters back to $\{\mathbf{B}_{k,\text{dq}}^{(t+1)},\mathbf{A}_{k,\text{dq}}^{(t+1)}\}$ and aggregates them using our proposed Adaptive Rank-1 Matrix-Level Aggregation method. This approach aggregates only the updated and important rank-1 matrices, avoiding information dilution and accelerating convergence compared to traditional zero-padding methods. These innovations enable the HAFLQ framework to achieve efficient communication, scalability, and high performance in federated learning.
II-C Wireless Communication Model
II-C1 Uplink Communication
In our HAFLQ framework, each communication round is of fixed duration $T$ and each client $k$ is allocated a limited bandwidth $B_{k}$. Consequently, the number of bits that a client can transmit in a given communication round is constrained. To characterize the wireless uplink between a client and the base station, we adopt a channel model that accounts for path loss, shadowing, and small-scale fading.
OFDMA-based Multi-User Access
When multiple clients are scheduled to transmit concurrently, an orthogonal frequency division multiple access (OFDMA) scheme is adopted to avoid inter-client interference. We let $\mathcal{C}$ denote the set of available subchannels, where each subchannel has identical bandwidth $B_{0}$. Then, a binary channel assignment vector $\bm{\rho}_{k}=[\rho_{k,1},\rho_{k,2},\ldots,\rho_{k,|\mathcal{C}|}]$ can be defined for client $k$, where $\rho_{k,c}=1$ indicates that subchannel $c$ is allocated to client $k$. The constraints
$\sum_{c\in\mathcal{C}}\rho_{k,c}$ $=1,\quad\forall k\in\mathcal{K},$ (4)
$\sum_{k\in\mathcal{K}}\rho_{k,c}$ $\leq 1,\quad\forall c\in\mathcal{C},$ (5)
ensure that each client is assigned exactly one subchannel and each subchannel is shared by at most one client in a given round. Since all subchannels have equal bandwidth $B_{0}$, the bandwidth allocated to client $k$ is $B_{k}=B_{0}$ for all $k\in\mathcal{K}$.
Channel Gain and Path Loss
We consider a common urban scenario where client $k$ has no line-of-sight (NLoS) to the base station[21] and the path loss is modeled as
$PL_{k}=32.4+20\log_{10}(f)+30\log_{10}(d_{k}),$ (6)
where $f$ is the carrier frequency and $d_{k}$ is the distance between client $k$ and the base station.
Incorporating shadowing and small-scale fading effects, the overall channel gain $h_{k}$ is given by
$h_{k}=10^{-PL_{k}/10}\,\psi_{k}\,\chi_{k},$ (7)
where the large-scale shadowing effect is modeled as
$\psi_{k}=10^{\xi_{k}/10},\quad\xi_{k}\sim\mathcal{N}(0,\sigma^{2}),$ (8)
and the small-scale Rayleigh fading component is
$\chi_{k}=|g|^{2},\quad\text{with }g\sim\mathcal{CN}(0,1).$ (9)
Here, $\mathcal{N}(0,\sigma^{2})$ denotes a real-valued Gaussian distribution with zero mean and variance $\sigma^{2}$, and $\mathcal{CN}(0,1)$ denotes a circularly symmetric complex Gaussian distribution with zero mean and unit variance.
Signal-to-Noise Ratio and Data Rate
Assuming that each client transmits with power $P_{k}$, the Signal-to-Noise Ratio (SNR) at the base station for client $k$ is given by
$\gamma_{k}=\frac{P_{k}\,h_{k}}{N_{0}^{\text{bs}}\,B_{k}},$ (10)
where $N_{0}^{\text{bs}}$ is the noise power spectral density (PSD) at the base station side. Under the Shannon capacity formulation, the achievable uplink data rate[22] of client $k$ is
$R_{k}=B_{k}\log_{2}\left(1+\gamma_{k}\right).$ (11)
Thus, during the uplink phase of a communication round, which lasts for $T_{\text{ul}}$, the total number of bits that client $k$ can transmit is
$M_{k}=T_{\text{ul}}R_{k}.$ (12)
Parameter Transmission
In each round, client $k$ transmits its updated LoRA parameters, which together occupy $S_{k}$ bits. Since the available bit budget per round is limited by (12), we require
$S_{k}\leq M_{k},$ (13)
for successful transmission.
II-C2 Downlink Communication
In the proposed HAFLQ framework, we assume that the base station broadcasts the global LoRA parameters $\Theta_{\text{LoRA}}^{(t)}=\{\mathbf{B}_{g}^{(t)},\mathbf{A}_{g}^{(t)}\}$ to all clients in full precision using shared channels. Given the low-rank nature of LoRA parameters, their dimensions are significantly smaller compared to the full model parameters, resulting in low communication overhead. Additionally, the high transmit power of the base station ensures that these parameters can be reliably delivered to all clients, even under varying channel conditions.
III Salience-Driven Adaptive LLM Quantization
In wireless FL systems, clients often have varying computational capabilities due to differences in hardware resources such as GPU memory and processing power. Deploying LLMs across such heterogeneous clients requires computation resource optimization, particularly GPU memory and computation.
To address this, we propose an adaptive quantization approach for LLMs that customizes the quantization level based on each client’s computational capability. Our method leverages a salience metric to assess the importance of transformer blocks, enabling selective quantization. Important transformer blocks, identified as having higher salience, are preserved in full precision to maintain model performance, while less important blocks are quantized to lower precision. The proportion of blocks quantized depends on the GPU memory available to the client. Clients with limited GPU memory quantize a larger proportion of transformer blocks to fit within their memory constraints, while clients with greater GPU memory retain more blocks in full precision to maximize model performance.
III-A Salience-Driven LLM Transformer Block Importance Evaluation
In our approach, the pre-trained LLM is composed of $L$ transformer blocks that are stacked sequentially. Each transformer block $l$ has parameters $\mathbf{W}_{\text{TB}_{l}}\in\mathbb{R}^{d\times d}$. To evaluate the importance of each transformer block, we adopt a salience-based metric inspired by SparseGPT [23], which quantifies the contribution of each weight element to the model’s output error. Specifically, for the weight matrix $\mathbf{W}_{\text{TB}_{l}}$ in a given transformer block, we compute the salience value $\delta_{i,j}$ for each element $w_{i,j}$ using
$\delta_{i,j}=\frac{w_{i,j}^{2}}{[\mathbf{H}^{-1}]_{j,j}}$ (14)
where $[\mathbf{H}^{-1}]_{j,j}$ is the $j$th diagonal element of inverse Hessian matrix $\mathbf{H}^{-1}$. The Hessian matrix $\mathbf{H}\in\mathbb{R}^{d\times d}$ is computed as the average outer product of input activation vectors $\mathbf{x}_{i}$ in the local dataset $\mathcal{D}_{k}$
$\mathbf{H}=\frac{1}{D_{k}}\sum_{i=1}^{D_{k}}\mathbf{x}_{i}\mathbf{x}_{i}^{T},$ (15)
where $\mathbf{x}_{i}$ is the input activation vector for the $i$th sample in the dataset $\mathcal{D}_{k}$. To efficiently compute the diagonal elements of $\mathbf{H}^{-1}$, we leverage Cholesky decomposition [24], which avoids explicitly inverting $\mathbf{H}$ and allows for efficient computation of the required diagonal elements.
Next, we calculate the total salience score for each transformer block $\mathbf{W}_{\text{TB}_{l}}$ by summing the salience values $\delta_{i,j}$ of all its weight elements
$R_{\text{TB}_{l}}=\sum_{i=1}^{d}\sum_{j=1}^{d}\delta_{i,j},$ (16)
the salience score $R_{\text{TB}_{l}}$ represents the overall impact of the transformer block $l$ on the model’s output error. Based on these salience scores, we rank all transformer blocks in descending order of importance.
This salience-driven evaluation provides the foundation for our quantization strategy. During quantization, transformer blocks with lower salience scores are prioritized for quantization, while those with higher salience scores are preserved in full precision to minimize the impact on model performance. The proportion of blocks to be quantized is dynamically adjusted based on the computational capacity and GPU memory size of the target device: for devices with higher computational power and larger GPU memory, only a small number of blocks are quantized, preserving most blocks in full precision. Conversely, for devices with lower computational power and smaller GPU memory, a larger proportion of blocks are quantized to meet resource constraints.
III-B Block-wise Quantization for LLMs
To efficiently quantify LLMs while minimizing the loss of accuracy, we utilize block-wise quantization [25, 26]. By dividing the large weight matrices into smaller blocks and quantizing each block independently, this method allows the quantization process to better adapt to the local distribution of weights.
III-B1 Divide the Weight Matrix
Given a weight matrix $\mathbf{W}_{\text{TB}_{l}}\in\mathbb{R}^{d\times d}$ from transformer block $l$, which is selected for quantization, the matrix is divided into several contiguous blocks along the row dimension. Specifically, assuming a block size of $m$, the matrix is partitioned into $p=d/m$ blocks: $\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{p}$, where each block $\mathbf{W}_{i}\in\mathbb{R}^{m\times d}$ contains $m$ consecutive rows of the original matrix.
III-B2 Quantize the Block
For each block $\mathbf{W}_{i}$, a scale factor $s_{i}$ is computed based on the range of values within the block using
$s_{i}=\frac{\max(\mathbf{W}_{i})-\min(\mathbf{W}_{i})}{2^{q}-1},$ (17)
where $q$ denotes the number of quantization bits. The quantized weights for block $\mathbf{W}_{i}$ are then obtained as
$\mathbf{W}_{i,\text{q}}=\left\lfloor\frac{\mathbf{W}_{i}}{s_{i}}\right\rceil,$ (18)
[…]
The quantized weights are stored in a compressed format in the device’s GPU memory, to ensure efficient access during computation. This storage format significantly reduces memory usage compared to full-precision weights, enabling the deployment of large models on resource-constrained devices.
III-B3 Dequantize the Block
The dequantization process occurs dynamically during the forward pass, both in training and inference. Specifically, when a quantized weight block $\mathbf{W}_{i,\text{q}}$ is needed for computation, it is dequantized on-the-fly using the corresponding scale $s_{i}$ as
$\mathbf{W}_{i}=s_{i}\cdot\mathbf{W}_{i,\text{q}}.$ (19)
This approach ensures that the computational overhead of dequantization is minimized in scenarios where only the required blocks are dequantized on-the-fly during computation, rather than the entire weight matrix.
IV Importance-based Parameter Truncation and Freezing Schemes